OpenStudy (anonymous):

tan(arctanx + arctan 1/4)=7/11, find x

OpenStudy (anonymous):

You should get the inverse of it

OpenStudy (anonymous):

trololololol

OpenStudy (anonymous):

hahahaha!

OpenStudy (anonymous):

i already did that and i just want to make sure that my answer is correct :))

OpenStudy (anonymous):

nakita nga lang kita eh

OpenStudy (anonymous):

ang susungit naman ng mga tao dito

OpenStudy (anonymous):

kala mo lang un.. may mabait dito :))

OpenStudy (anonymous):

just use tan double angle formula.....

hartnn (hartnn):

arctanA + arctan B = arctan((A+B)/(1-AB))

OpenStudy (anonymous):

wait.. i'm confused @_@

OpenStudy (anonymous):

i'm confused too

OpenStudy (anonymous):

mee too

hartnn (hartnn):

arctanx + arctan 1/4 = arctan((x+1/4)/(1-x/4)) tan(arctanx + arctan 1/4) = (x+1/4)/(1-x/4) =7/11

hartnn (hartnn):

now can u find x ? (x+1/4)/(1-x/4) =7/11